OpenStudy (anonymous):

The 2000 census found a U.S. population of about 281 million. Write an equation for the U.S. population that assumes exponential growth at 0.7% per year. Use the equation to predict the U.S. population in 2100. equation: y = P*(1+r)x P is the initial value (value of y when x = 0), r is the percent change (written as a decimal), and x is the input variable and y is the output variable.

OpenStudy (anonymous):

281*(1.007)^100 = 564.49 Million You wrote the equation wron, it should be y = P*(1+r)^x

OpenStudy (campbell_st):

the formula is \[y = P \times(1 + r)^x\] then using the \[y = 281000000(1 = 0.07)^{100}\]available information

OpenStudy (campbell_st):

x is the difference between the initial population year 2000 and the year 2100

OpenStudy (anonymous):

yes i got that answer (564,486,260) but when I try to show it on excel it doesnt fit with my data. ill post a screenshot of my table.

OpenStudy (anonymous):

notice how the last number does not fit with the rest of the data above, there's different differences if that makes sense.

OpenStudy (anonymous):

it is .7% not 7%

OpenStudy (anonymous):

\[y = 281000000(1.007)^{100}\]

OpenStudy (anonymous):

.7% = .007, not .07